A person with normal eyesight uses a converging lens of focal length 10 cm. He holds the lens close to his eye to finish an engraving job. At what distance from the job he must hold the lens?

-Answer=7.1; why?
Answer #1

we see anything most clearly wehen its 25cm from our eye. so feye=25cm(feye is variable but its most relaxed at this point)

now if he uses another converging lens of focal length 10 cm then the new focal length will be

1/fnew=1/feye+1/flens

or,1/fnew=1/25+1/10=7/50

so, fnew=7.1cm so he must put the object 7.1 cm from his eye